FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF ACETYLENE-1,1-DIFLUOROETHYLENE

\underline{HELEN O. LEUNG}, Department of Chemistry, Amherst College, P.O. Box 5000, Amherst, MA 01002-5000.

The rotational spectra of four isotopomers of HCCH-CH_{2}CF_{2} (the most abundant isotopomer, HCCD-CH_{2}CF_{2}, DCCD-CH_{2}CF_{2}, and H^{13}C^{13}CH-CH_{2}CF_{2}) have been collected in the 6-18 GHz region with a pulsed molecular beam, Fourier transform microwave spectrometer. The nuclear quadrupole hyperfine structure due to the deuterium nuclei in two of the isotopomers is observed. The spectroscopic constants are consistent with a planar structure in which the primary interaction is between H in HCCH and one of the F atoms in 1,1-difluoroethylene. There is also a secondary interaction between the acetylenic bond and the H atom cis to the F atom involved in the hydrogen bond with HCCH.